A projectile is fired into the air from the top of a 200-m cliff above a valley. Its initial velocity is 60 m/s at 60° above the horizontal. (a) Where does the projectile land? (b) How long does the ball in air? (c) At what times velocity is 40m/s? I.
Answer #1

rounc -200-CoSin60)七- ( 94 ) t 3)S 969.8t 700 (51-96-98 t 25 .45 = 51.96-98ト 七= 2.6sec At time t -2